# AutoSTF: Decoupled Neural Architecture Search for Cost-Effective Automated Spatio-Temporal Forecasting

### v1

@@ Abstract.
Spatio-temporal forecasting is a critical component of various smart city applications, such as transportation optimization, energy management, and socio-economic analysis. Recently, several automated spatio-temporal forecasting methods have been proposed to automatically search the optimal neural network architecture for capturing complex spatio-temporal dependencies. However, the existing automated approaches suffer from expensive neural architecture search overhead, which hinders their practical use and the further exploration of diverse spatio-temporal operators in a finer granularity. In this paper, we propose AutoSTF, a decoupled automatic neural architecture search framework for cost-effective automated spatio-temporal forecasting. From the efficiency perspective, we first decouple the mixed search space into temporal space and spatial space and respectively devise representation compression and parameter-sharing schemes to mitigate the parameter explosion. The decoupled spatio-temporal search not only expedites the model optimization process but also leaves new room for more effective spatio-temporal dependency modeling. From the effectiveness perspective, we propose a multi-patch transfer module to jointly capture multi-granularity temporal dependencies and extend the spatial search space to enable finer-grained layer-wise spatial dependency search. Extensive experiments on eight datasets demonstrate the superiority of AutoSTF in terms of both accuracy and efficiency. Specifically, our proposed method achieves up to 13.48$\times$ speed-up compared to state-of-the-art automatic spatio-temporal forecasting methods while maintaining the best forecasting accuracy.
@@ 1. Introduction
Spatio-temporal forecasting is the process of predicting future states that depend on both spatial and temporal contexts (Jiang et al., 2023; Lin et al., 2022; Han et al., 2024). For example, in transportation systems, the future traffic flow and speed can be estimated by simultaneously learning the spatio-temporal dependencies of historical traffic conditions generated by geo-distributed roadside sensors (Liu et al., 2020; Fang et al., 2021; Zhang et al., 2024). Accurate spatio-temporal forecasting plays a pivotal role in various smart city applications, such as human mobility modeling (Li et al., 2022a; Cirstea et al., 2022; He et al., 2018), demand-supply rebalancing (Zhang et al., 2020b; Ling et al., 2023), and urban anomaly event detection (Tedjopurnomo et al., 2020; LIU et al., 2022; Han et al., 2023; Liu et al., 2022).
Figure 1. The training time and forecasting accuracy (mean squared error) comparison of manually-designed (MTGNN, GWN, DCRNN) and automated (AutoSTG, AutoCTS) spatio-temporal forecasting models on the METR-LA dataset. Our proposed method (AutoSTF) achieves the best forecasting accuracy while taking much less training time compared with all existing automated models.
@@ 2. Preliminaries
[…]
AutoSTF Framework.
@@ 3. The AutoSTF Framework
In this section, we provide a comprehensive overview of our proposed model. Additionally, we give a detailed analysis in Appendix A to explain why the model is designed in this way.
@@ 4.1. Experimental Settings
Datasets. We performed experiments on eight benchmark datasets to evaluate both multi-step and single-step forecasting. For multi-step forecasting, we utilized traffic speed datasets (METR-LA and PEMS-BAY) provided by Li et al. (Li et al., 2017) and traffic flow datasets (PEMS03, PEMS04, PEMS07, and PEMS08) released by Song et al. (Song et al., 2020). As for single-step forecasting, we employed solar-energy and electricity datasets made available by Lai et al. (Lai et al., 2018). We provide the detailed statistics of the datasets in Appendix B.1.
Evaluation Metrics. We followed previous works (Wu et al., 2021; Pan et al., 2021; Wu et al., 2023) to evaluate the model’s performance in multi- and single-step forecasting. Specifically, Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE) were utilized to assess the accuracy of multi-step forecasting. For single-step forecasting, we utilized Root Relative Squared Error (RRSE) and Empirical Correlation Coefficient (CORR) as evaluation metrics. More details are stated in Appendix B.3.
Baselines. We compare AutoSTF with 11 manual-designed methods, including DCRNN (Li et al., 2017), STGCN (Yu et al., 2018), Graph WaveNet (Wu et al., 2019), AGCRN (Bai et al., 2020), MTGNN (Wu et al., 2020), STID (Shao et al., 2022b), STHODE (Yao et al., 2023), DeepSTUQ (Qian et al., 2023), STJGCN (Zheng et al., 2023), STAEformer (Liu et al., 2023) and LightCTS (Lai et al., 2023), and 3 automated models including AutoSTG (Pan et al., 2021), AutoCTS+ (Wu et al., 2023) and AutoCTS (Wu et al., 2021). We provide the detailed description of baselines in Appendix B.2.
Detailed information on hyper-parameters and implementation specifics can be found in Appendix B.4.
4.2. Overall Results of AutoSTF
Firstly, we present the main experiments for both multi-step (refer to Table 1 and 2) and single-step (see Table 3) spatio-temporal forecasting. We adopt the convention of highlighting the best accuracy in bold and underlining the second best accuracy.
Multi-step forecasting. As demonstrated in Tables 1 and 2, AutoSTF outperforms all baseline models on all multi-step forecasting tasks across all evaluation metrics. In summary, the results demonstrate that AutoSTF can produce highly competitive architectures adept at capturing spatio-temporal dependencies, leading to enhanced forecasting accuracy.
@@ 4.3. Efficiency Study
4.4. Ablation Study
5. Related Work
The manually-designed models. Existing manually-designed models often propose complex algorithms to model the spatial-temporal dependencies using domain-specific knowledge (Liu et al., 2023; Zheng et al., 2023; Deng et al., 2022; Yao et al., 2023). In particular, many works (Deng et al., 2024; Pan et al., 2019; Yu et al., 2018; Li et al., 2023b; Shao et al., 2022a; Jiang et al., 2023) involve Spatial-Temporal Graph Neural Networks (STGNN) to explore the temporal dependencies within time series and spatial correlations across different locations for spatio-temporal forecasting. For temporal modeling, techniques such as Gated Recurrent Units (GRU), Long Short-Term Memory (LSTM) (Sutskever et al., 2014), and Temporal Convolutional Networks (TCN) (Yu and Koltun, 2016) are commonly employed to extract information from time series data. In addition, various Graph Neural Network approaches have been proposed for modeling spatial dependencies using different adjacency matrices, including predefined and adaptive matrices (Wu et al., 2019; Li et al., 2017). Graph WaveNet (Wu et al., 2019), as proposed by Wu et al., utilizes both predefined and adaptive matrices to capture spatio-temporal dependencies. Furthermore, several studies have proposed methods to model dynamic spatial dependencies (Han et al., 2021; Shao et al., 2022c; Zhang et al., 2022; Ye et al., 2022). These methods primarily focus on learning the hidden relationships between nodes using dynamic node features, which are generally characterized by a combination of real-time traffic attributes. For example, ESG (Ye et al., 2022) presents a flexible graph neural network that models multi-scale time series interactions to simultaneously capture pairwise correlations and temporal dependencies. However, manually-designed models necessitate substantial domain-specific expertise, as well as extensive parameter adjustments.
@@ 6. Conclusion
In this paper, we propose AutoSTF, a novel decoupled neural architecture search framework for cost-effective automated spatio-temporal forecasting. By decoupling the mixed search space into temporal and spatial space, we design representation compression and parameter-sharing schemes to mitigate the parameter explosion in our framework. The decoupled spatio-temporal search not only expedites the model optimization process but also leaves new room for more effective spatio-temporal dependency modeling. In addition, in the multi-patch transfer module, we aim to jointly capture multi-granularity temporal dependencies and extend the spatial search space to enable finer-grained layer-wise spatial dependency search. In spatial search, we design layer-wise message-passing spatial operators to capture the spatio-temporal dependencies by automatically selecting the optimal adjacency matrix in different layers, thereby enhancing forecasting accuracy. Comprehensive experiments conducted on eight datasets highlight the superior performance of AutoSTF in terms of both accuracy and efficiency. In future work, focusing on improving the search algorithm, expanding the search space, and scaling the model to handle large datasets presents a valuable and meaningful challenge.
@@ A.2. The Efficiency Bottleneck
FF Distance-based Matrix DCRNN (Li et al., 2017), GrapWaveNet (Wu et al., 2019), DGCN (Pan et al., 2019), STGNN (Wang et al., 2020), GMAN (Zheng et al., 2020), ST-GRAT (Park et al., 2020), STDN (Yu et al., 2018)
@@ A.3. The Effectiveness Bottleneck
In addition, existing automated methods have employed the same message-passing scheme across different GNN layers, which fails to consider the unique spatial dependencies present in different multi-hop neighbors. Recent advances have proposed diverse adjacency matrices for spatial dependency modeling and demonstrated the effectiveness of different adjacency matrices in different contexts  (Jin et al., 2022). As summarized in Table 7, we categorize the commonly used adjacency matrices in three categories (Li et al., 2023a): (1) the fully fixed FF matrix that is the predefined adjacency matrix and fixed in both the training and inferencing phase; (2) the semi-adaptive SA matrix that the parameters of the adjacency matrix are learnable with the training phase, and fixed in the inferencing phase; (3) the fully adaptive FA matrix that was learned from data-driven methods and adaptive in both the training and inferencing phase. In fact, each adjacency matrix category can be regarded as a new class of spatial operators. It is intuitive to incorporate varying adjacency matrix operators in different layers so as to encompass more spatial semantics for more effective forecasting. In this work, we construct an extended spatial search space that includes the distance matrix, adaptive matrix, and attention matrix operators. Based on the new spatial space, we propose a more flexible message-passing scheme to automatically select the most suitable adjacency matrix in different layers for better forecasting accuracy.
@@ B.4. Implementation Details
The experiments were conducted on three different server configurations. The first configuration consisted of three Linux Centos servers, each equipped with 4 RTX 3090 GPUs. The second configuration included one Linux Ubuntu server with 2 V100 GPUs, and the third configuration included one Linux Ubuntu server with 2 A40 GPUs.
Following the previous works (Wu et al., 2021; Pan et al., 2021; Liu et al., 2019), we use MAE as the loss function to train the model and use Adam with a learning rate of 0.001 for both the architecture parameters $\Theta$ and the network weights $\omega$. In addition, we use a weight decay of 0.0001 as the optimizer, and the batch size is set to 64. In the temporal and spatial DAG, we set the number of nodes as 4.

### v2

@@ Abstract.
Spatio-temporal forecasting is a critical component of various smart city applications, such as transportation optimization, energy management, and socio-economic analysis. Recently, several automated spatio-temporal forecasting methods have been proposed to automatically search the optimal neural network architecture for capturing complex spatio-temporal dependencies. However, the existing automated approaches suffer from expensive neural architecture search overhead, which hinders their practical use and the further exploration of diverse spatio-temporal operators in a finer granularity. In this paper, we propose AutoSTF, a decoupled automatic neural architecture search framework for cost-effective automated spatio-temporal forecasting. From the efficiency perspective, we first decouple the mixed search space into temporal space and spatial space and respectively devise representation compression and parameter-sharing schemes to mitigate the parameter explosion. The decoupled spatio-temporal search not only expedites the model optimization process but also leaves new room for more effective spatio-temporal dependency modeling. From the effectiveness perspective, we propose a multi-patch transfer module to jointly capture multi-granularity temporal dependencies and extend the spatial search space to enable finer-grained layer-wise spatial dependency search. Extensive experiments on eight datasets demonstrate the superiority of AutoSTF in terms of both accuracy and efficiency. Specifically, our proposed method achieves up to 13.48$\times$ speed-up compared to state-of-the-art automatic spatio-temporal forecasting methods while maintaining the best forecasting accuracy. The source code and data are available at https://github.com/usail-hkust/AutoSTF.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709323isbn: 979-8-4007-1245-6/25/08ccs: Information systems Spatial-temporal systemsccs: Computing methodologies Neural networks
@@ 1. Introduction
Spatio-temporal forecasting is the process of predicting future states that depend on both spatial and temporal contexts (Wang et al., 2020a; Jiang et al., 2023; Lin et al., 2022). For example, in transportation systems, the future traffic flow and speed can be estimated by simultaneously learning the spatio-temporal dependencies of historical traffic conditions generated by geo-distributed roadside sensors (Zhang and Yan, 2022; Liu et al., 2020; Fang et al., 2021). Accurate spatio-temporal forecasting plays a pivotal role in various smart city applications, such as human mobility modeling (Li et al., 2022a; Cirstea et al., 2022; Han et al., 2021a), demand-supply rebalancing (Zhang et al., 2020b; Ling et al., 2023), and urban anomaly event detection (Tedjopurnomo et al., 2020; LIU et al., 2022).
Figure 1. The training time and forecasting accuracy (mean squared error) comparison of manually-designed and automated spatio-temporal forecasting models on the METR-LA dataset. Our proposed method (AutoSTF) achieves the best forecasting accuracy while taking much less training time compared with all existing automated models.
@@ 2. Preliminaries
Table 1. The temporal and spatial operators of search space.
Type Operators Formulary
$\mathcal{T}$ $GDCC$ $GDCC(X)=XW_{1}\odot\sigma(XW_{2}))$
$Informer$ $Informer(X)=Softmax(\frac{(XW_{Q})(XW_{K})^{T}}{\sqrt{D^{\prime}}}(XW_{V}))$
$\mathcal{S}$ $GNN_{fixed}$ $GNN_{fixed}(X)=\sum_{k=0}^{K}\mathbf{A}_{dis}^{k}XW_{fixed}$
$GNN_{adap}$ $GNN_{adap}(X)=\sum_{k=0}^{K}\mathbf{A}_{adap}^{k}XW_{adap}$
$GNN_{att}$ $GNN_{att}(X)=(ReLU(\mathbf{A}_{att}XW_{i}^{V}\cdot W_{1}+b_{1}))\cdot W_{2}+b_{2}$ $s.t.\mathbf{A}_{att}=softmax(\frac{XW_{t}^{Q}\cdot(XW_{i}^{K})^{T}}{\sqrt{D^{\prime}}})$
@@ 3. The AutoSTF Framework
In this section, we provide a comprehensive overview of our proposed model. Additionally, we give a detailed analysis in Appendix A to explain why the model is designed in this way. We list all the operators of temporal search and spatial search in Table 1.
@@ 4.1. Experimental Settings
In this subsection, we describe the datasets used, the metrics employed in our experiments, and the baseline comparisons we have conducted. Detailed information on hyper-parameters and implementation specifics can be found in Appendix B.4.
Datasets. We performed experiments on eight benchmark datasets to evaluate both multi-step and single-step forecasting. For multi-step forecasting, we utilized traffic speed datasets (METR-LA and PEMS-BAY) provided by Li et al. (Li et al., 2017) and traffic flow datasets (PEMS03, PEMS04, PEMS07, and PEMS08) released by Song et al. (Song et al., 2020). As for single-step forecasting, we employed solar-energy and electricity datasets made available by Lai et al. (Lai et al., 2018). These datasets can be categorized as traffic speed, traffic flow, solar-energy, and electricity. We provide the detailed statistics of the datasets in Appendix B.1.
Evaluation Metrics. We followed previous works (Wu et al., 2021; Pan et al., 2021; Wu et al., 2023) to evaluate the model’s performance in multi- and single-step forecasting. Specifically, Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE) were utilized to assess the accuracy of multi-step forecasting. For single-step forecasting, we utilized Root Relative Squared Error (RRSE) and Empirical Correlation Coefficient (CORR) as evaluation metrics. Lower values of MAE, RMSE, MAPE, and RRSE are indicative of higher accuracy, while larger CORR values denote higher accuracy. More details are stated in Appendix B.3.
@@
Baselines. We compare AutoSTF with 14 manual-designed methods, including DCRNN (Li et al., 2017), STGCN (Yu et al., 2018), Graph WaveNet (Wu et al., 2019), AGCRN (Bai et al., 2020), MTGNN (Wu et al., 2020), STID (Shao et al., 2022b), STHODE (Yao et al., 2023), DeepSTUQ (Qian et al., 2023), STJGCN (Zheng et al., 2023), STAEformer (Liu et al., 2023) and LightCTS (Lai et al., 2023), and 3 automated models including AutoSTG (Pan et al., 2021), AutoCTS+ (Wu et al., 2023) and AutoCTS (Wu et al., 2021). We provide the detailed description of baselines in Appendix B.2.
4.2. Overall Results of AutoSTF
Firstly, we present the main experiments for both multi-step (refer to Table 2 and 3) and single-step (see Table 4) spatio-temporal forecasting. We adopt the convention of highlighting the best accuracy in bold and underlining the second best accuracy.
Multi-step forecasting. As demonstrated in Tables 2 and 3, AutoSTF outperforms all baseline models on all multi-step forecasting tasks across all evaluation metrics. In summary, the results demonstrate that AutoSTF can produce highly competitive architectures adept at capturing spatio-temporal dependencies, leading to enhanced forecasting accuracy.
@@ 4.3. Efficiency Study
4.4. Ablation Study
5. Related Work
The manually-designed models. Existing manually-designed models often propose complex algorithms to model the spatial-temporal dependencies using domain-specific knowledge (Liu et al., 2023; Zheng et al., 2023; Qian et al., 2023; Yao et al., 2023). In particular, many works (Li et al., 2017; Pan et al., 2019; Yu et al., 2018; Li et al., 2023b; Shao et al., 2022a; Jiang et al., 2023) involve Spatial-Temporal Graph Neural Networks (STGNN) to explore the temporal dependencies within time series and spatial correlations across different locations for spatio-temporal forecasting. For temporal modeling, techniques such as Gated Recurrent Units (GRU), Long Short-Term Memory (LSTM) (Sutskever et al., 2014), and Temporal Convolutional Networks (TCN) (Yu and Koltun, 2016) are commonly employed to extract information from time series data. In addition, various Graph Neural Network approaches have been proposed for modeling spatial dependencies using different adjacency matrices, including predefined and adaptive matrices (Wu et al., 2019; Li et al., 2017). Graph WaveNet (Wu et al., 2019), as proposed by Wu et al., utilizes both predefined and adaptive matrices to capture spatio-temporal dependencies. Furthermore, several studies have proposed methods to model dynamic spatial dependencies (Han et al., 2021b; Shao et al., 2022c; Zhang et al., 2022; Ye et al., 2022). These methods primarily focus on learning the hidden relationships between nodes using dynamic node features, which are generally characterized by a combination of real-time traffic attributes. For example, ESG (Ye et al., 2022) presents a flexible graph neural network that models multi-scale time series interactions to simultaneously capture pairwise correlations and temporal dependencies. However, manually-designed models necessitate substantial domain-specific expertise, as well as extensive parameter adjustments.
@@ 6. Conclusion
In this paper, we propose AutoSTF, a novel decoupled neural architecture search framework for cost-effective automated spatio-temporal forecasting. By decoupling the mixed search space into temporal and spatial space, we design representation compression and parameter-sharing schemes to mitigate the parameter explosion in our framework. The decoupled spatio-temporal search not only expedites the model optimization process but also leaves new room for more effective spatio-temporal dependency modeling. In addition, in the multi-patch transfer module, we aim to jointly capture multi-granularity temporal dependencies and extend the spatial search space to enable finer-grained layer-wise spatial dependency search. In spatial search, we design layer-wise message-passing spatial operators to capture the spatio-temporal dependencies by automatically selecting the optimal adjacency matrix in different layers, thereby enhancing forecasting accuracy. Comprehensive experiments conducted on eight datasets highlight the superior performance of AutoSTF in terms of both accuracy and efficiency. In future work, focusing on improving the search algorithm, expanding the search space, and scaling the model to handle large datasets presents a valuable and meaningful challenge. Additionally, developing a method to automatically determine the optimal number of patches for different datasets is another meaningful task.
@@ A.2. The Efficiency Bottleneck
FF Distance-based Matrix DCRNN (Li et al., 2017), GrapWaveNet (Wu et al., 2019), DGCN (Pan et al., 2019), STGNN (Wang et al., 2020b), GMAN (Zheng et al., 2020), ST-GRAT (Park et al., 2020), STDN (Yu et al., 2018)
@@ A.3. The Effectiveness Bottleneck
In addition, existing automated methods have employed the same message-passing scheme across different GNN layers, which fails to consider the unique spatial dependencies present in different multi-hop neighbors. Recent advances have proposed diverse adjacency matrices for spatial dependency modeling and demonstrated the effectiveness of different adjacency matrices in different contexts  (Jin et al., 2022). As summarized in Table 8, we categorize the commonly used adjacency matrices in three categories (Li et al., 2023a): (1) the fully fixed (FF) matrix that is the predefined adjacency matrix and fixed in both the training and inferencing phase; (2) the semi-adaptive (SA) matrix that the parameters of the adjacency matrix are learnable with the training phase, and fixed in the inferencing phase; (3) the fully adaptive (FA) matrix that was learned from data-driven methods and adaptive in both the training and inferencing phase. In fact, each adjacency matrix category can be regarded as a new class of spatial operators. It is intuitive to incorporate varying adjacency matrix operators in different layers so as to encompass more spatial semantics for more effective forecasting. In this work, we construct an extended spatial search space that includes the distance matrix, adaptive matrix, and attention matrix operators. Based on the new spatial space, we propose a more flexible message-passing scheme to automatically select the most suitable adjacency matrix in different layers for better forecasting accuracy.
@@ B.4. Implementation Details
The experiments were conducted on three different server configurations. The first configuration consisted of three Linux Centos servers, each equipped with 4 RTX 3090 GPUs. The second configuration included one Linux Ubuntu server with 2 V100 GPUs, and the third configuration included one Linux Ubuntu server with 2 A800 GPUs.
Following the previous works (Wu et al., 2021; Pan et al., 2021; Liu et al., 2019), we use MAE as the loss function to train the model and use Adam with a learning rate of 0.001 for both the architecture parameters $\Theta$ and the network weights $\omega$. In addition, we use a weight decay of 0.0001 as the optimizer, and the batch size is set to 64. In the temporal and spatial DAG, we set the number of nodes as 4. In order to ensure reproducibility, we have provided the detailed settings for each dataset in our source code.